Hello. Can you help me with this equation?
2x-3/(x+4)=x-3/x-4
I have tried to cross multiply by multiplying 2x-3(x+4) and (x-3)(x-4) which gave me 2x^2+8x-3x-12 = x^2-4x-3x+12. From here I simplified the equations so they looked like this:
2x^2+5x-12 = x^2-x+12 but here's where I stumble. The answer in the book is
(0,18) but I'm not getting this.

Answer by Boreal(15235)

Couple of things:
First, the answers given (0,18) do not work for this equation.
You get -(3/4)=(-3/-4). Eighteen doesn't work, either.
Cross multiplying gives you (2x-3)(x-4) =(x-3)(x+4), which is a little different.
Still, that ends up with x^2-12x+24=0
(1/2) 12 +/- sqrt (144-96)=
(1/2) 12 +/- 4 sqrt (3)
6+2 sqrt 3
6-2 sqrt 3

just using sqrt (3)=1.732
9.464 would be a solution
18.928-3/13.464 on the left
6.464/5.464 on the right
1.1830=1.1830
Check to be sure parentheses I had were what the problem was, the problem is copied correctly, and the answer in the back of the book is truly what they say it is. Clearly, zero does not work